What is the temperature (in °C) of an 11.8 L sample of carbon monoxide, CO, at 749 torr if it occupies 12.5 L at 55.3°C and 749 torr?
Answer #1

V1/T1 = V2 / T2 (ideal gas equation is PV= nRT)

11.8 / T1 = 12.5 / 55.3

11.8 / T1 = 0.226

11.8 / 0.226 = T1

T1 = 52.21 °C